# A Survey on Unlearning in Large Language Models

### v1

@@ Abstract.
The advancement of Large Language Models (LLMs) has revolutionized natural language processing, yet their training on massive corpora poses significant risks, including the memorization of sensitive personal data, copyrighted material, and knowledge that could facilitate malicious activities. To mitigate these issues and align with legal and ethical standards such as the “right to be forgotten”, machine unlearning has emerged as a critical technique to selectively erase specific knowledge from LLMs without compromising their overall performance. This survey provides a systematic review of over 180 papers on LLM unlearning published since 2021, focusing exclusively on large-scale generative models. Distinct from prior surveys, we introduce novel taxonomies for both unlearning methods and evaluations. We clearly categorize methods into training-time, post-training, and inference-time based on the training stage at which unlearning is applied. For evaluations, we not only systematically compile existing datasets and metrics but also critically analyze their advantages, disadvantages, and applicability, providing practical guidance to the research community. In addition, we discuss key challenges and promising future research directions. Our comprehensive overview aims to inform and guide the ongoing development of secure and reliable LLMs.
@@
copyright: noneccs: Information systems Language modelsccs: Security and privacy
@@ 1. Introduction
Large Language Models (LLMs) have significantly transformed research paradigms in natural language processing while enabling a diverse array of practical applications. These capabilities arise from training on extensive textual corpora, which allows the models to internalize and encode substantial knowledge within their parameters. However, this capacity also introduces critical risks. For instance, personally identifiable information memorized during training can be extracted through privacy attacks, raising concerns under data protection regulations such as the “right to be forgotten” (Voigt and Von dem Bussche, 2017; Regulation, 2016). Similarly, unauthorized use of copyrighted materials in training data can expose model providers to legal challenges (Wei et al., 2025a). Moreover, LLMs can internalize knowledge that facilitates malicious activities (Li et al., 2024; Lang et al., 2025), and jailbreak attacks can elicit the generation of harmful or illegal content. In light of these concerns, selectively erasing specific knowledge from LLMs has emerged as a necessary step toward enhancing their security, reliability, and regulatory compliance.
@@
One potential solution is to retrain LLMs from scratch after removing problematic data. However, this approach is computationally expensive and impractical for large-scale models. Machine unlearning (Cao and Yang, 2015) offers a more efficient alternative, which aims to develop algorithms to selectively remove the influence of specific training data while preserving the overall performance of the model on retained data. In the context of LLMs, the distinctive autoregressive next-token prediction mechanism (Yao et al., 2024a) has motivated extensive research into unlearning methods specifically designed for these models. This survey narrows its focus to address unlearning techniques tailored for large-scale generative language models, which are predominantly used for generative tasks rather than classification. [^1] Some early LLM unlearning works also considered classification tasks in natural language processing (Pawelczyk et al., 2024; Bhaila et al., 2025; Chen and Yang, 2023), but they are not the focus of this survey. By systematically reviewing more than 180 papers published since 2021222Some articles were retrieved from public repositories such as https://github.com/chrisliu298/awesome-llm-unlearning., this survey aims to provide a comprehensive overview of the definition, methods, evaluations, challenges and future directions in LLM unlearning.
Several existing surveys touch upon LLM unlearning, some of which adopt a broader scope or concentrate on specialized aspects (Cooper et al., 2024; Zhou et al., 2024; Liu et al., 2024c; Xu, 2024; Barez et al., 2025; Zhang et al., 2025b; Si et al., 2023; Qu et al., 2025). Compared to surveys that also focus specifically on LLM unlearning (Geng et al., 2025a; Blanco-Justicia et al., 2025), this work offers a more systematic and comprehensive perspective, with several distinctive contributions in the following paragraph. A detailed comparison is summarized in Table 1.
@@
(1) A novel taxonomy of unlearning methods. We categorize unlearning approaches based on the training stage at which they are applied: training time, post-training, and inference time. This taxonomy offers a clearer organizational structure compared to alternative classifications based on parameter selection, since some full-parameter methods can also be applied to part of the parameters or by incorporating LoRA adapters to apply to extra parameters. (2) Multidimensional analysis of evaluations. Instead of merely enumerating existing datasets and metrics, we provide a multidimensional analysis for both datasets and metrics. For datasets, through a comparison from the perspectives of task format, content, and experimental paradigms, we evaluate the characteristics of 18 existing benchmarks, offering actionable guidance for researchers. For metrics, from the goal of LLM unlearning, we analyze knowledge memorization metrics and their applicability, along with commonly used metrics for model utility, robustness, and efficiency. (3) Discussion of Challenges and Future Directions. We provide an in-depth discussion of current challenges in LLM unlearning and outline prospective research directions. These contributions aim to accelerate progress in the emerging field of LLM unlearning, ultimately contributing to safer and more responsible AI systems.
      Current Survey Feature Comparison
      Cooper et al. (2024); Zhou et al. (2024); Liu et al. (2024c) Generative AI Limited attention to generative LLMs.
      Xu (2024) Traditional and large-scale models
      Barez et al. (2025) Open problems Specific aspects of LLM unlearning.
      Zhang et al. (2025b) Legal perspectives
      Si et al. (2023); Qu et al. (2025) Classification-oriented settings
      Geng et al. (2025a); Blanco-Justicia et al. (2025) LLM unlearning (1) Classify by parameters modified.
(2) Enumerate datasets and metrics.
Table 1. A comparison of our work with existing surveys on LLM unlearning. Most surveys focus on a larger or smaller scope than ours. Two surveys similar to ours focusing on LLM Unlearning adopt different classification criteria and lack detailed analysis on datasets and metrics.
@@ 2.1. Machine Unlearning in LLMs
Within the standard framework of machine unlearning, we consider a dataset $\mathcal{D}$ and an original model $\mathcal{M}$, parameterized by $\theta$, trained on $\mathcal{D}$. The subset of training data targeted for removal is denoted as the unlearn set $\mathcal{D}_{u}\subset\mathcal{D}$, while the remainder constitutes the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{u}$. The objective of machine unlearning is to design an algorithm $\mathcal{U}$ that takes as input the original model $\mathcal{M}$ and the relevant data, and outputs an unlearned model $\mathcal{M}_{u}$. This model $\mathcal{M}_{u}$ is intended to approximate the behavior of a retrained model $\mathcal{M}_{r}$, which is trained exclusively on the retain set $\mathcal{D}_{r}$. An illustration of the unlearning process is provided in Figure 1. In the context of LLM unlearning, we provide specific explanations from two aspects: (1) different types of unlearning request and (2) the goal of unlearning.
@@ 2.1.1. Type of Unlearning Request
Anallise Ivory was born on November 8, 1990, and her Social Security Number is 900-55-1236. (Ramakrishna et al., 2025)
@@
“There’s more in the frying pan,” said Aunt Petunia, turning eyes on her massive son. (Shi et al., 2025)
@@
Samples: Stephen King is a world-renowned American author of horror, suspense, supernatural fiction, … (Jin et al., 2024)
@@ Abstract
Samples: Adversaries may use brute force techniques to gain access to … (Lang et al., 2025)
@@
The predominant form of unlearning request operates at the sample level, requiring models to forget specific text sequences that contain sensitive information, thereby mitigating privacy (Tian et al., 2024; Ramakrishna et al., 2025), copyright (Eldan and Russinovich, 2024; Shi et al., 2025), or safety risks (Li et al., 2024). Examples of different samples are shown in Figure 2. These sequences may consist of free-form text or structured question–answer pairs, as outlined in Table 4.
Beyond isolated samples, a growing number of work addresses entity-level unlearning, which aims at removing all knowledge associated with a particular entity. Entities may be concrete (e.g., individuals, books) (Jin et al., 2024; Choi et al., 2025) or abstract (e.g., biases, capabilities) (Lang et al., 2025), as depicted in Figure 2. Usually, this task is reduced to sample-level unlearning by constructing a corresponding unlearn set with samples related to the target entity. Compared to sample-level unlearning, it requires not only erasing memorized content but also managing inter-entity correlations, rendering it significantly more challenging.
@@ 2.2.1. Memorization and Data Extraction
As formalized by the goal of unlearning in Section 2.1.2, the conceptualization of memorization directly shapes the objectives of unlearning, while the methodology for memorization detection provides an essential diagnostic tool for evaluating unlearning efficacy. There exist multiple definitions of LLM memorization, such as formulations based on counterfactual memorization (Zhang et al., 2023a) and tuple completion (Meng et al., 2022). Among these, extractable memorization (Carlini et al., 2021) is the most prevalent, conceptualizing memorization as content that the model can reproduce under specific prompting conditions. This definition originally involved identifying a precise input prefix to induce the model to output the memorized content, and has evolved into a diverse class of data extraction attacks, employing various input strategies and detection mechanisms (Shin et al., 2020; Zou et al., 2023). Consequently, data extraction attacks serve a dual role: they constitute a critical tool for evaluating unlearning, particularly for assessing the knowledge memorization, while unlearning itself functions as a defensive measure to purge hazardous knowledge and thereby mitigate the risks posed by malicious data extraction attempts.
@@ 2.2.2. Knowledge Updating
Knowledge editing and updating are essential for maintaining the long-term efficacy of large language models (LLMs), as they enable the correction of inaccuracies and the integration of new knowledge without requiring full model retraining. LLM unlearning can be viewed as a promising strategy within this domain, with research advancing in two main directions: some studies develop robust, conflict-free parameter update algorithms to facilitate reliable knowledge modification (Jung et al., 2025; Ni et al., 2024; Sun et al., 2024), while others apply unlearning techniques to domain-specific contexts (Yang and Lo, 2024; Fore et al., 2024). Another widely adopted paradigm is model editing, which focuses on local, targeted modifications to specific factual knowledge while preserving the model’s general capabilities and avoiding catastrophic forgetting. A key distinction between model editing and unlearning lies in their objectives: model editing operates with a predefined target knowledge state, whereas unlearning aims to remove or suppress information without necessarily replacing it. Nevertheless, mechanistic insights from model editing techniques, such as knowledge neurons and locate-then-edit approaches (Meng et al., 2022), can inform the design of more precise and interpretable unlearning methods.
@@ 2.2.3. Alignment
Alignment seeks to ensure that LLMs behave in accordance with human values and intentions. This objective is dual in nature, involving the guide of models toward generating helpful responses (positive) and preventing them from producing undesirable outputs (negative). A widely adopted approach for positive guidance is reinforcement learning from human feedback (RLHF), which steers models toward desirable behaviors through iterative reward-based optimization (Bai et al., 2022; Ouyang et al., 2022). Complementarily, LLM unlearning has emerged as a critical technique for negative alignment, systematically removing undesirable knowledge or capabilities from models. For example, it has been applied to mitigate social biases (Yu et al., 2023; Dige et al., 2024), eliminate unauthorized content to protect copyright (Wei et al., 2025a), and reduce the risk of leaking sensitive information (Rashid et al., 2025; Geng et al., 2025b). Together, these methods form a cohesive alignment framework that addresses both the promotion of beneficial behaviors and the suppression of harmful ones.
@@ 3. Existing unlearning methods
Output Refer to caption Trained Model TrainingInference SISA-based 3.1 Training-time Input Output Token Embedding Token Logit Instruction Sample Decode Offset 3.3 Inference-time Fine-tuning Others Objectives Parameters Text Distribution Activation Multiple objective Full Partial Extra Data-based Data-free Arithmetic operations SAE-based PEFT Subspace 3.2 Post-training
Figure 3. Framework of unlearning methods. In typical LLM usage scenarios, a model is first trained on specific datasets, and then is used for inference to generate outputs. The unlearning method can be applied to the training process, the trained model, or the inference stage, corresponding to training-time unlearning (Section 3.1, post-training unlearning (Section 3.2) and inference-time unlearning (Section 3.3).
In typical LLM usage scenarios, a model is first trained on specific datasets from draft or a pretrained base model, and then is used for inference to generate output in some tasks. As illustrated in Figure 3, the unlearning method can be applied to the training process, the trained model, or the inference stage, corresponding to training-time unlearning (Section 3.1), post-training unlearning (Section 3.2) and inference-time unlearning (Section 3.3). In short, Training-Time Unlearning requires adjusting the training process to facilitate unlearning, which is mainly based on SISA training paradigms. Post-Training Unlearning involves altering the trained model, mainly through fine-tuning towards multiple objectives on selected parameters. Inference-time Unlearning aims to achieve unlearning via input or output adjustments, rather than modifying the model parameters.
@@ 3.1. Training-Time Unlearning
As noted in Section 2.1, the ideal outcome of unlearning is to obtain a retrained model $\mathcal{M}_{r}$. However, full retraining starting from $\mathcal{M}_{b}$ is computationally prohibitive and impractical in real-world scenarios. To alleviate this burden, training-time unlearning techniques, exemplified by SISA (Bourtoule et al., 2021), introduce novel training frameworks that initiate retraining from intermediate states. These methods partition the dataset into multiple subsets and store corresponding checkpoints trained on different subsets. By constraining the influence of data points, retraining can start from specific checkpoints unaffected by the unlearn set, thereby accelerating the unlearning process.
Given the considerable storage and computational overhead associated with maintaining numerous model copies, Bannihatti Kumar et al. (2023) and Chowdhury et al. (2025) integrate supplementary trainable components, applying the SISA principle to fine-tune and preserve only the newly introduced parameters. This strategy substantially reduces the number of parameters requiring updates. Beyond efficiency and performance considerations, Kadhe et al. (2023) examine fairness concerns in SISA-based frameworks and propose FairSISA, which integrates three post-processing bias mitigation techniques.

### v2

@@ Abstract.
Large Language Models (LLMs) demonstrate remarkable capabilities, but their training on massive corpora poses significant risks from memorized sensitive information. To mitigate these issues and align with legal standards, unlearning has emerged as a critical technique to selectively erase specific knowledge from LLMs without compromising their overall performance. This survey provides a systematic review of over 180 papers on LLM unlearning published since 2021. First, it introduces a novel taxonomy that categorizes unlearning methods based on the phase in the LLM pipeline of the intervention. This framework further distinguishes between parameter modification and parameter selection strategies, thus enabling deeper insights and more informed comparative analysis. Second, it offers a multidimensional analysis of evaluation paradigms. For datasets, we compare 18 existing benchmarks from the perspectives of task format, content, and experimental paradigms to offer actionable guidance. For metrics, we move beyond mere enumeration by dividing knowledge memorization metrics into 10 categories to analyze their advantages and applicability, while also reviewing metrics for model utility, robustness, and efficiency. By discussing current challenges and future directions, this survey aims to advance the field of LLM unlearning and the development of secure AI systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Language modelsccs: Security and privacy
@@ 1. Introduction
Large Language Models (LLMs) have significantly transformed research paradigms in natural language processing while enabling a diverse array of practical applications. These capabilities arise from training on extensive textual corpora, which allows the models to encode substantial knowledge within their parameters. However, this capacity also introduces critical risks. For instance, personally identifiable information memorized during training can be extracted through privacy attacks, raising concerns under data protection regulations such as “right to be forgotten” (voigt2017eu; regulation2016regulation). Similarly, unauthorized use of copyright materials in training data can expose model providers to legal challenges (Wei2024Evaluating). Moreover, LLMs can internalize knowledge that facilitates malicious activities (Li2024WMDP; lang2025beyond), and jailbreak attacks can elicit the generation of harmful or illegal content. In light of these concerns, selectively erasing specific knowledge from LLMs has emerged as a necessary step toward enhancing their security, reliability, and regulatory compliance.
@@
One potential solution is to retrain LLMs from scratch after removing problematic data. However, this approach is computationally expensive and impractical for large-scale models. Machine unlearning (Cao2015Making) offers a more efficient alternative, which aims to develop algorithms to selectively remove the influence of specific training data while preserving the overall performance of the model on retained data. In the context of LLMs, the distinctive autoregressive next-token prediction mechanism (yao2024machine) has motivated extensive research into unlearning methods specifically designed for these models. This survey narrows its focus to address unlearning techniques tailored for large-scale generative language models, which are predominantly used for generative tasks rather than classification. [^1] Some LLM unlearning works also considered classification tasks in natural language processing (Pawelczyk2024Context; Bhaila2024Soft; Chen2023Unlearn), but they are not the focus of this survey.
Several existing surveys touch upon LLM unlearning. Most of them either adopt a broader scope (cooper2024machine; zhou2024limitations; liu2024machine; xu2024machine), concentrate on specialized aspects (barez2025open; zhang2024right; Si2023Knowledge; qu2024frontier), or lack coverage of extensive research published after October 2024 (blanco2025digital; le-khac2025survey). To address these gaps, we provide a comprehensive and up-to-date overview by systematically reviewing more than 180 papers published since 2021.222Some articles were retrieved from public repositories such as https://github.com/chrisliu298/awesome-llm-unlearning. In contrast to other recent surveys on LLM unlearning (ren2025sok; geng2025comprehensive), our work introduces a novel taxonomy and delivers an in-depth analysis, with specific contributions outlined below.
(1) A novel taxonomy of unlearning methods (Section 3). We categorize unlearning approaches based on their execution timing: training time, post-training, and inference time. Compared to alternative taxonomies based on unlearning objectives or intentions, our framework provides a clearer organizational structure with two advantages. First, it distinguishes between parameter modification and parameter selection strategies, offering a flexible conceptual space for their integration, and thereby enabling deeper insights. For example, while RMU (Li2024WMDP) is grouped under categories such as “localized parameter modification” in the prior survey, our taxonomy separately analyzes its loss design in the SFT section (3.2.1) and its parameter selection mechanism in the parameter localization section (3.2.3). Second, our taxonomy is usage-oriented, enabling the informed selection and comparative analysis of unlearning methods according to specific operational scenarios.
@@
(2) Multidimensional analysis of evaluations (Section 4). Instead of merely enumerating existing datasets and metrics, we provide a multidimensional analysis for both datasets and metrics. For datasets, through a comparison from the perspectives of task format, content, and experimental paradigms, we evaluate the characteristics of 18 existing benchmarks, offering actionable guidance for researchers. For metrics, from the goal of LLM unlearning, we divide knowledge memorization metrics into 10 categories to analyze their advantages and applicability, along with commonly used metrics for model utility, robustness, and efficiency.
(3) Discussion of Challenges and Future Directions (Section 5). We provide an in-depth discussion of current challenges in LLM unlearning, including the lack of a strict and consistent definition, the variation of impacts across languages and data, and the difficulties of real-world implementation. Furthermore, we outline prospective research directions, aiming to advance the field of LLM unlearning and contribute to more responsible AI systems.
@@ 2.1. Machine Unlearning in LLMs
Within the standard framework of machine unlearning, we consider a dataset $\mathcal{D}$ and an original model $\mathcal{M}$, parameterized by $\theta$, trained on $\mathcal{D}$. The subset of training data targeted for removal is denoted as the unlearn set $\mathcal{D}_{u}\subset\mathcal{D}$, while the remainder constitutes the retain set $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{u}$. The objective of machine unlearning is to design an algorithm $\mathcal{U}$ that takes the original model $\mathcal{M}$ and the relevant data as input and produces an unlearned model $\mathcal{M}_{u}$. This model $\mathcal{M}_{u}$ is intended to approximate the behavior of a retrained model $\mathcal{M}_{r}$, which is trained exclusively on the retain set $\mathcal{D}_{r}$. An illustration of the unlearning process is provided in Figure 1. In the context of LLM unlearning, we provide specific explanations from two aspects: (1) different types of unlearning request and (2) the goal of unlearning.
@@ 2.1.1. Type of Unlearning Request
Anallise Ivory was born on November 8, 1990, and her Social Security Number is 900-55-1236. (ramakrishna2025lume)
@@
“There’s more in the frying pan,” said Aunt Petunia, turning eyes on her massive son. (Shi2024MUSE)
@@
Samples: Stephen King is a world-renowned American author of horror, suspense, supernatural fiction, … (Jin2024RWKU)
@@ Abstract
Samples: Adversaries may use brute force techniques to gain access to … (lang2025beyond)
@@
The predominant form of unlearning request operates at the sample level, requiring models to forget specific text sequences that contain sensitive information, thereby mitigating privacy (Tian2024Forget; ramakrishna2025lume), copyright (Eldan2023Whos; Shi2024MUSE), or safety risks (Li2024WMDP). Examples of different samples are shown in Figure 2. These sequences may consist of free-form text or structured question–answer pairs, as outlined in Table 3.
Beyond isolated samples, a growing number of work addresses entity-level unlearning, which aims at removing all knowledge associated with a particular entity. Entities may be concrete (e.g., individuals, books) (Jin2024RWKU; choi2025opt-out) or abstract (e.g., biases, capabilities) (lang2025beyond), as depicted in Figure 2. Usually, this task is reduced to sample-level unlearning by constructing a corresponding unlearn set with samples related to the target entity. Compared to sample-level unlearning, it requires not only erasing memorized content but also managing inter-entity correlations, rendering it significantly more challenging.
@@ 2.2.1. Memorization and Data Extraction
As formalized by the goal of unlearning in Section 2.1.2, the conceptualization of memorization directly shapes the objectives of unlearning, while the methodology for memorization detection provides an essential diagnostic tool for evaluating unlearning efficacy. There exist multiple definitions of LLM memorization, such as formulations based on counterfactual memorization (zhang2023counterfactual) and tuple completion (meng2022locating). Among these, extractable memorization (carlini2021extracting) is the most prevalent, conceptualizing memorization as content that the model can reproduce under specific prompting conditions. This definition originally involved identifying a precise input prefix to induce the model to output the memorized content, and has evolved into a diverse class of data extraction attacks, employing various input strategies and detection mechanisms (shin2020autoprompt; zou2023universal). Consequently, data extraction attacks serve a dual role: they constitute a critical tool for evaluating unlearning, particularly for assessing the knowledge memorization, while unlearning itself functions as a defensive measure to purge hazardous knowledge and thereby mitigate the risks posed by malicious data extraction attempts.
@@ 2.2.2. Knowledge Updating
Knowledge editing and updating are essential for maintaining the long-term efficacy of large language models (LLMs), as they enable the correction of inaccuracies and the integration of new knowledge without requiring full model retraining. LLM unlearning can be viewed as a promising strategy within this domain, with research advancing in two main directions: some studies develop robust, conflict-free parameter update algorithms to facilitate reliable knowledge modification (jung2025come; ni2024forgetting; sun2024learning), while others apply unlearning techniques to domain-specific contexts (yang2024hotfixing; fore2024unlearning). Another widely adopted paradigm is model editing, which focuses on local, targeted modifications to specific factual knowledge while preserving the model’s general capabilities and avoiding catastrophic forgetting. A key distinction between model editing and unlearning lies in their objectives: model editing operates with a predefined target knowledge state, whereas unlearning aims to remove or suppress information without necessarily replacing it. Nevertheless, mechanistic insights from model editing techniques, such as knowledge neurons and locate-then-edit approaches (meng2022locating), can inform the design of more precise and interpretable unlearning methods.
@@ 2.2.3. Alignment
Alignment seeks to ensure that LLMs behave in accordance with human values and intentions. This objective is dual in nature, involving the guide of models toward generating helpful responses (positive) and preventing them from producing undesirable outputs (negative). A widely adopted approach for positive guidance is reinforcement learning from human feedback (RLHF), which steers models toward desirable behaviors through iterative reward-based optimization (bai2022training; ouyang2022training). Complementarily, LLM unlearning has emerged as a critical technique for negative alignment, systematically removing undesirable knowledge or capabilities from models. For example, it has been applied to mitigate social biases (Yu2023Unlearning; Dige2024Mitigating), eliminate unauthorized content to protect copyright (Wei2024Evaluating), and reduce the risk of leaking sensitive information (Rashid2024Forget; geng2025mitigating). Together, these methods form a cohesive alignment framework that addresses both the promotion of beneficial behaviors and the suppression of harmful ones.
@@ 3. Existing unlearning methods
Output Refer to caption Trained Model TrainingInference SISA-based 3.1 Training-time Input Output Token Embedding Token Logit Instruction Sample Decode Offset 3.3 Inference-time How to update Which parameter 3.2.1 SFT 3.2.2 RL Text Distribution Activation Multiple objective Full 3.2.3 Partial 3.2.4 Extra Data-based Data-free 3.2.5 Composite Arithmetic SAE-based 3.2 Post-training
Figure 3. Framework of unlearning methods. In typical LLM usage scenarios, a model is first trained on specific datasets, and then is used for inference to generate outputs. The unlearning method can be applied to the training process, the trained model, or the inference stage, corresponding to training-time unlearning (Section 3.1, post-training unlearning (Section 3.2) and inference-time unlearning (Section 2).
In typical LLM usage scenarios, a model is first trained on specific datasets from draft or a pretrained base model, and then is used for inference to generate output in some tasks. As illustrated in Figure 3, the unlearning method can be applied to the training process, the trained model, or the inference stage, corresponding to training-time unlearning (Section 3.1), post-training unlearning (Section 3.2) and inference-time unlearning (Section 2).
In short, Training-Time Unlearning requires adjusting the training process to facilitate unlearning, which is mainly based on SISA training paradigms. Post-Training Unlearning involves altering the trained model, mainly through supervised fine-tuning (SFT) or reinforcement training (RL) on selected parameters. Inference-time Unlearning aims to achieve unlearning via input or output adjustments, rather than modifying the model parameters.
@@ 3.1. Training-Time Unlearning
As noted in Section 2.1, the ideal outcome of unlearning is to obtain a retrained model $\mathcal{M}_{r}$. However, full retraining starting from $\mathcal{M}_{b}$ is computationally prohibitive and impractical in real-world scenarios. To alleviate this burden, training-time unlearning techniques, exemplified by SISA (Bourtoule2021Machine), introduce novel training frameworks that initiate retraining from intermediate states. These methods partition the dataset into multiple subsets and store corresponding checkpoints trained on different subsets. By constraining the influence of data points, retraining can start from specific checkpoints unaffected by the unlearn set, thereby accelerating the unlearning process.
Given the considerable storage and computational overhead associated with maintaining numerous model copies, BannihattiKumar2023Privacy and Chowdhury2024Scalable integrate supplementary trainable components, applying the SISA principle to fine-tune and preserve only the newly introduced parameters. This strategy substantially reduces the number of parameters requiring updates. Beyond efficiency and performance considerations, kadhe2023fairsisa examine fairness concerns in SISA-based frameworks and propose FairSISA, which integrates three post-processing bias mitigation techniques.